A student conducts a simple random sample of students from her high school and finds that 32 out of 100 students in her sample regularly walk to school. Give a point estimate for the proportion of all students at her high school who regularly walk to school.
Answer #1

Solution :

Given that,

n = 100

x = 32

Point estimate = sample proportion = \hat p = x / n = 32/ 100=0.32

answer = 0.32